OpenStudy (anonymous):

I'm on lecture 3, Electric Flux & Gauss's Law, and Prof Lewin drops in epsilon 0. Based on the previous two lectures, I have no idea what epsilon 0 represents. Anybody know? Reply: I just searched wikipedia and saw that it's the electric constant.

OpenStudy (fifciol):

according to Coulomb's law: F~q1q2/r^2 F=kq1q2/r^2 k=1/4pi epsilon0 epsilon0 is just a number, some constant. such that the law is true

OpenStudy (anonymous):

got it. cheers

OpenStudy (fifciol):

yw:)

OpenStudy (anonymous):

epsilon 0 represents the permittivity of free space(vaccum) and has a value of 8.85*10^(-12)

OpenStudy (anonymous):

It represents the ability of vacuum to store electrical energy in a material or the extent of polarization of a material in vacuum both in presence of an electric field